OpenStudy (anonymous):

find the distance between the points (1,4) and (4,8)

OpenStudy (anonymous):

\[\sqrt{(1-4)^2+(4-8)^2}=5\]

OpenStudy (anonymous):

\[\sqrt{(4-1)^{2}+(8-4)^{2}}\] \[=\sqrt{3^{2}+4^{2}}=\sqrt{9+16}=\sqrt{25}=5\]